To devise effective marketing strategies it is helpful to know the characteristics of your customers. A study compared the demographic characteristics of people who use the Internet for travel arrangements. Of 932 Internet users, 543 had completed college. (a) Find a 95% confidence interval for the proportion of people who use the Internet for travel arrangements, and have a college degree. (b) Can you conclude from this study that the majority of people who use the Internet for travel arrangements own a college degree? Justify your answer using the confidence interval computed in (a).
Answer #1

Here the no.of internet users (n) = 932

No.of people use internet for travel (x) = 543

sample proportion = x/n = 543 / 932 = 0.58262

Here use the Z- tables to find the critical value at 95% confidence